OpenStudy (anonymous):

Find two numbers with a sum of 20 and a difference of 14

OpenStudy (mertsj):

x+y=20 x-y=14 Solve the system

pooja195 (pooja195):

try to add the numbers in each answer and see which one gets you 20, try to subtract the numbers in each answer to see what gives you 14.

pooja195 (pooja195):

one number can be 2 for example

OpenStudy (anonymous):

A. 4 and 18 B. –3 and –17 C. 2 and 16 D. 3 and 17

pooja195 (pooja195):

ok so wat do u think?

OpenStudy (anonymous):

D. 3 AND 17

pooja195 (pooja195):

correct good job!! :)

OpenStudy (anonymous):

HAVE A PHONE

OpenStudy (anonymous):

x+y=3 y=2x-15

pooja195 (pooja195):

yes good